OpenStudy (anonymous):

I need help solving the initial value problem: dy/dx - (y/x) = xe^(x), y(1) = e - 1

OpenStudy (anonymous):

Once again, we have an equation in the form: \[\frac{dy}{dx}+p(x)y=q(x)\] In this case: \[\eqalign{ &p(x)=-\frac{1}{x} \\ &q(x)=x(e)^x \\ }\]

OpenStudy (anonymous):

So we can isolate for y: \[y=\frac{\int{e^{\int{-\frac{1}{x}dx}}x(e)^x}+c_1}{e^{\int{-\frac{1}{x}dx}}}\]

OpenStudy (anonymous):

And we arrive to the general equation: \[y=x(e^x+c_1)\] Now, we can sub in \(y(1)=e^1-1\) And derive the specific equation which would be the answer, you tell me if you want it

OpenStudy (anonymous):

\[I.F=e ^{\int\limits \frac{ -1 }{x}dx}=e ^{-\ln x}=e ^{\ln x ^{-1}}=x ^{-1}=\frac{ 1 }{ x }\] \[C.S is y *\frac{ 1 }{ x }=\int\limits x e ^{x}*\frac{ 1 }{ x }dx+c\] \[\frac{ y }{x }=e ^{x}+c\]

OpenStudy (anonymous):

when x=1, y=e-1 use it and find the value of c

OpenStudy (anonymous):

Keith, for the last equation above would it be - C_1

OpenStudy (anonymous):

?

OpenStudy (anonymous):

If I'm to use y(1) = e^1 - 1, aren't I simply plugging x = 1 into the general formula so that it equals the equation: y(1) = e^1 -1. I'm just a little confused when you said derive the specific equation

OpenStudy (anonymous):

Ohh well when they say \(y(1)=e^1-1\), That means that there exists the point: \[\eqalign{ &y=e^1-1 \\ &x=1 \\ }\] So we can plug this in like so: \[y=x(e^x+c_1)\] \[e^1-1=1(e^1+c_1)\] And solve for c, then plug it back into \[y=x(e^x+c_1)\]

OpenStudy (anonymous):

Anyways, Patrick, I got to go, I will leave the calculation of \(c_1\) to you and good luck haha If im online, and you would like me to take a look at any question, don't hesitate to give me a shout m'kay?

OpenStudy (anonymous):

Ok I see. I actually got the answer correct as its in the book. Thank you for your help

OpenStudy (anonymous):

Anytime!